# Imitative Membership Inference Attack

### v1

@@ Appendix B Additional Experimental Results
Figure 19: The ROC curves of attack results on ResNet50 models trained on four image datasets in the non-adaptive setting.
@@ B.1 Effectiveness of Imitative Training
Likelihood-Based Analysis. Likelihood is a widely used metric for assessing the fitness of a given value to a distribution. Here, we use it to quantify how closely the target model’s true confidence scores align with the distributions learned by imitative models and shadow models. Specifically, for both members and non-members, we compute the likelihood of the target model’s confidence score under the estimated Gaussian distribution of imitative and shadow models. We then define the Averaged Likelihood Ratio as the mean likelihood of imitative models divided by that of shadow models, averaged over all members and non-members. A ratio greater than one indicates that the imitative model’s distribution more accurately approximates the target model’s behavior than the shadow model’s distribution. As shown in Table˜20, this ratio is consistently and substantially above one across datasets, which provides strong quantitative evidence that imitative training creates more faithful and accurate mimics of the target model’s behavior.
@@ B.2 Additional Ablation Studies
Attack with Distribution Shift. In realistic attack scenarios, the adversary’s data is often not perfectly aligned with the target’s data. We follow prior work [3, 19] and conduct the following experiments to evaluate $\mathsf{IMIA}$ in such cases:
@@
  • $\mathbb{D}_{\text{target}}\neq\mathbb{D}_{\text{attack}}$. The target model is trained on CIFAR-10, while the imitative models use the ImageNet [8] portion of the CINIC-10 [7], creating a distribution shift between the target’s data and the adversary’s data.
@@
Table 21: Incorporating the CMIA framework [9] for $\mathsf{IMIA}$.
@@
TPR @ 0.001%FPR TPR @ 0.1%FPR Balanced Accuracy
@@
1 1.33% 2.33% 2.35% 3.61% 54.25% 61.30%
3 1.42% 2.55% 2.54% 3.71% 54.38% 61.32%
6 1.45% 2.59% 2.62% 3.73% 54.38% 61.32%
9 1.45% 2.60% 2.65% 3.73% 54.40% 61.33%
B.3 Incorporating with CMIA
CMIA [9] is a newly proposed attack-agnostic framework that enhances the performance of shadow-based MIAs by iteratively training models on carefully selected datasets. We incorporate this framework with $\mathsf{IMIA}$ and report results in Table˜21. As shown, the performance improves with more cascading iterations, particularly during the initial few stages, which is consistent with observations from the original paper. Nevertheless, CMIA is applicable only in the adaptive setting and incurs additional computational overhead, which limits its practicality.
B.4 Experiments on Other Models in the Non-Adaptive Settings
We present the experimental results on the remaining three model architectures in the non-adaptive setting. Table˜22 shows the results for VGG16, Table˜23 presents the results for DenseNet121, and Table˜24 displays the results for MobileNetV2. We also present the ROC curves of non-adaptive attacks on four datasets, as shown in Figure˜19.
@@ B.4 Experiments on Other Models in the Non-Adaptive Settings
Method TPR @ 0.001% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
Method TPR @ 0.001% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
Method TPR @ 0.001% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
B.5 Experiments on Other Models in the Adaptive Settings
We present the results on the remaining three model architectures in the adaptive setting.  Table˜25 shows the results for VGG16, Table˜26 presents the results for DenseNet121, and Table˜27 displays the results for MobileNetV2.
@@ B.5 Experiments on Other Models in the Adaptive Settings
Method TPR @ 0.001% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
Method TPR @ 0.001% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
Method TPR @ 0.001% FPR TPR @ 0.1% FPR Balanced Accuracy

### v2

@@ Appendix B Additional Experimental Results
Figure 19: The ROC curves of attack results on ResNet models trained on four image datasets in the non-adaptive setting.
@@ B.1 Effectiveness of Imitative Training
Likelihood-Based Analysis. Likelihood is a widely used metric for assessing the fitness of a given value to a distribution. Here, we use it to quantify how closely the target model’s true confidence scores align with the distributions learned by imitative models and shadow models. Specifically, for both members and non-members, we compute the likelihood of the target model’s confidence score under the estimated Gaussian distribution of imitative and shadow models. We then define the Averaged Likelihood Ratio as the mean likelihood of imitative models divided by that of shadow models, averaged over all members and non-members. A ratio greater than one indicates that the imitative model’s distribution more accurately approximates the target model’s behavior than the shadow model’s distribution. As shown in Table˜19, this ratio is consistently above one across all datasets, providing strong quantitative evidence that imitative training creates more faithful and accurate mimics of the target model’s behavior.
@@ B.2 Additional Ablation Studies
Table 20: Impact of query budget on attack performance (i.e., TPR @ 0%FPR). Querying the model with multiple augmented views enhances attack performance, yielding significant gains with just two queries.
Queries Non-Adaptive Setting Adaptive Setting
CIFAR-10 CIFAR-100 CIFAR-10 CIFAR-100
1 0.73% 1.07% 1.53% 5.54%
2 1.22% 1.89% 2.06% 7.69%
9 1.43% 2.04% 2.31% 8.42%
18 1.45% 2.10% 2.33% 8.52%
36 1.45% 2.11% 2.33% 8.55%
Impact of Query Budget. We evaluate the impact of varying query budgets on CIFAR-10 and CIFAR-100. As shown in Table˜20, augmenting random queries (e.g., horizontal flips and shifts by up to $\pm 4$ pixels) improves attack performance. Notably, significant gains occur with just two queries, and performance stabilizes around 18 queries. This observation aligns with findings in prior work [sp22lira].
@@
Attack with Distribution Shift. In realistic attack scenarios, the adversary’s data is often not perfectly aligned with the target’s data. We follow prior work [sp22lira, ccs24rapid] and conduct the following experiments to evaluate $\mathsf{IMIA}$ in such cases:
@@
  • $\mathbb{D}_{\text{target}}\neq\mathbb{D}_{\text{attack}}$. The target model is trained on CIFAR-10, while the imitative models use the ImageNet [cvpr09imagenet] portion of the CINIC-10 [arxiv18cinic], creating a distribution shift between the target’s data and the adversary’s data.
@@
Table 21: Comparison of attack performance (i.e., TPR @ 0% FPR) between LiRA and $\mathsf{IMIA}$ using CMIA [ndss26cpmia]. The runtime is measured (in hours) with 9 cascading iterations of both LiRA and $\mathsf{IMIA}$ using an A100 GPU.
@@
MNIST FMNIST CIFAR-10 CIFAR-100
LiRA $\mathsf{IMIA}$ LiRA $\mathsf{IMIA}$ LiRA $\mathsf{IMIA}$ LiRA $\mathsf{IMIA}$
1 0.80% 1.33% 3.85% 4.62% 1.30% 2.33% 5.41% 8.52%
3 1.05% 1.42% 4.29% 4.86% 2.05% 2.55% 7.49% 10.65%
6 1.46% 1.53% 4.72% 4.92% 2.26% 2.59% 9.30% 11.42%
9 1.53% 1.65% 4.90% 4.98% 2.53% 2.60% 10.62% 11.45%
Runtime (hrs) 128.8 5.8 129.8 5.9 1281.2 60.5 2562.3 120.9
B.3 Incorporating with CMIA
CMIA [ndss26cpmia] is a recently proposed attack-agnostic framework that enhances the performance of shadow-based MIAs by iteratively training models on carefully selected datasets. We incorporate this framework with $\mathsf{IMIA}$ and compare with LiRA. The result in Table˜21 shows that the performance improves with more cascading iterations, particularly during the initial few stages, which is consistent with observations from the original paper. We observe that CMIA provides a much larger boost to LiRA than to $\mathsf{IMIA}$; after 9 iterations, LiRA eventually reaches performance levels comparable to $\mathsf{IMIA}$. We think this is because CMIA trains more reliable shadow models by constructing datasets that resemble the target’s training data. This is similar to the goal of imitative training, where we aim to train target-informed imitative models. As a result, CMIA’s impact is more noticeable for target-agnostic attacks like LiRA. However, as shown in the table, applying CMIA with LiRA incurs significant computational overhead (over $20\times$ higher than $\mathsf{IMIA}$) and is only feasible in the adaptive setting, which limits its practicality.
B.4 Impact of Relaxing Assumptions
Our experiments adopt the standard assumptions used in the MIA literature. Here, we examine the impact of relaxing these assumptions to demonstrate the robustness of $\mathsf{IMIA}$.
Equal-size Assumption. While the standard evaluation constructs the query set with an equal number of members and non-members, our attack, like existing MIAs, does not rely on this assumption. First, the training of imitative models is independent of the query set’s composition; the adversary randomly selects instances to train imitative out models and selects pivot instances to train imitative in models regardless of the query set distribution. Second, our primary evaluation metric, TPR at 0% FPR, measures the highest true positive rate while permitting no false positives. This metric remains valid regardless of how the query set is balanced.
Table 22: Attack performance under different overlap ratios between the adversary’s dataset and the query set.
# Overlap
ratio
TPR @ 0%FPR TPR @ 0.1%FPR Balanced Accuracy
@@ B.4 Impact of Relaxing Assumptions
0 1.01% 1.45% 1.86% 3.42% 54.14% 61.08%
1/10 1.04% 1.44% 1.79% 3.46% 54.12% 61.04%
1/5 1.03% 1.43% 1.84% 3.48% 54.18% 61.10%
1/3 1.08% 1.57% 1.85% 3.52% 54.15% 61.20%
Disjoint Assumption. In the non-adaptive setting, prior work [sp22lira, ndss26cpmia] assumes that the adversary’s dataset and the query set are disjoint, i.e., $D_{\text{adv}}^{\text{non-adapt}}\cap D_{\text{query}}=\emptyset$. This assumption may not hold in realistic scenarios, where high-probability instances can appear in both datasets. $\mathsf{IMIA}$ handles this overlap straightforwardly. If a query instance $(x,y)$ appears in the adversary’s dataset $D_{\text{adv}}^{\text{non-adapt}}$, we simply discard the imitative out models trained with $(x,y)$ and use only those where $(x,y)$ was held out. In expectation, this requires twice as many imitative models to maintain the same number of imitative out models per query instance.
We use this approach to evaluate our attack under varying degrees of overlap between $D_{\text{adv}}^{\text{non-adapt}}$ and $D_{\text{query}}$. The results in Table˜22 show that attack performance remains stable across different overlap ratios. We also note that this strategy is highly conservative: when substantial overlap exists, one could leverage both the imitative in and imitative out behaviors for overlapped instances (same as the adaptive version of $\mathsf{IMIA}$) to achieve even stronger performance. Overall, these experiments demonstrate that our attack does not depend on the disjoint assumption and continues to perform reliably when it is relaxed.
B.5 Experiments on Other Models
We extend our evaluation to three model architectures: VGG16, DenseNet121, and MobileNetV2. For the non-adaptive setting, the detailed results are provided in Tables˜23, 24 and 25, with corresponding ROC curves shown in Figure˜19. Similarly, for the adaptive setting, the results are reported in Tables˜26, 27 and 28.
@@ B.5 Experiments on Other Models
Method TPR @ 0% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
Method TPR @ 0% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
Method TPR @ 0% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
Method TPR @ 0% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
Method TPR @ 0% FPR TPR @ 0.1% FPR Balanced Accuracy
@@
Method TPR @ 0% FPR TPR @ 0.1% FPR Balanced Accuracy